# PromptPex: Automatic Test Generation for Language Model Prompts

### v1

@@ 4.4. Benchmarks
We list the prompts with sources and their descriptions in Table 1. We selected a diverse set of prompts from publicly available sources, focusing on those that are within the scope of PromptPex. Currently, we support prompts that can accept only a single input. Although this single input can be interpreted by the model to be made up of multiple components, this differentiation is not explicitly present, for example, a single string as input representing a sentence and a word for the speech tag prompt. This limitation makes it unsuitable for prompts requiring multiple embedded inputs. We also only support prompts where the output is independent of the previous outputs, making prompts describing multi-turn conversations for tasks out of scope for PromptPex.
4.5. Evaluation Procedure
@@
We accessed gpt-4o and gpt-4o-mini through APIs. We kept the temperature 1.0 across all the requests. We used Ollama (ollama, ) for the local models, gemma2:9b, qwen2.5:3b, llama3.2:1b. We ran all the experiments and hosted the local models on a virtual machine (VM) hosted on Microsoft Azure. Our system runs on an AMD EPYC 7V13 processor with 256 GB of RAM, 2 TB of SSD, and an NVIDIA A100 GPU with 80 GB of dedicated memory. The VM is running Ubuntu 22.04.5 LTS and is allocated 24 out of 64 CPU cores. All the prompts are written and ran using GenAIScript (genaiscript, ) and Prompty (prompty, ). We ran each test once per prompt per model.
4.6. How useful are specifications for generating tests? (RQ1)
To test our hypothesis that having explicit specifications like OR and IS helps in generating better tests, we generated tests for the different prompts in our benchmark using PromptPex and the baseline. To compare the quality of test generation, we compare the percentage of tests not compliant with the prompt description for PromptPex and the baseline on multiple models. Higher percentage of non-compliance is better as it represents more test failures. Table 2 shows the percentage of test non-compliance for different prompts on each model. PromptPex on average generated 5.5% more tests for different models that were not compliant with description of the PUT from the benchmark.
@@ 4.6. How useful are specifications for generating tests? (RQ1)
Figure 10 shows the average non-compliance across the benchmarks for PromptPex and baseline for each of the different models. The more capable models (like gpt-4o-mini) have lower rates of non-compliance while the smaller models (like llama3.2:1b) have high rates. PromptPex beats baseline in every case but is particularly more effective for gpt-4o-mini, where the baseline model has a much harder time generating any tests that result in non-compliant output.
Prompts gpt-4o-mini gemma2:9b qwen2.5:3b llama3.2:1b
PPex BL PPex BL PPex BL PPex BL
speech-tag 0% 0% 5% 2% 2% 7% 90% 98%
text-to-p 15% 5% 23% 38% 72% 69% 95% 97%
shakespeare 2% 2% 5% 2% 5% 5% 7% 7%
sentence 12% 0% 4% 6% 21% 2% 31% 19%
extract-names 0% 0% 6% 16% 31% 22% 61% 46%
elements 31% 2% 44% 15% 62% 19% 54% 62%
classify 4% 0% 0% 0% 12% 0% 25% 25%
art-prompt 2% 0% 17% 12% 17% 7% 48% 45%
Average 8% 1% 13% 11% 28% 16% 51% 50%
Table 2. Test non-compliance results for the tests generated by PromptPex(PPex) vs the baseline (BL) on different models. More test non-compliance (winner shown in blue) represents more challenging tests and hence are better tests.
Prompts gpt-4o-mini gemma2:9b qwen2.5:3b llama3.2:1b
RL Inv RL Inv RL Inv RL Inv
speech-tag 0% 0% 0% 10% 5% 0% 95% 86%
text-to-p 14% 17% 14% 33% 62% 83% 90% 100%
shakespeare 0% 5% 0% 10% 0% 10% 0% 14%
sentence 17% 8% 0% 8% 25% 17% 38% 25%
extract-names 0% 0% 0% 11% 12% 28% 56% 67%
elements 33% 29% 17% 71% 58% 67% 50% 58%
classify 0% 8% 0% 0% 0% 25% 25% 25%
art-prompt 0% 5% 5% 29% 10% 24% 29% 67%
Average 8% 9% 4% 21% 24% 32% 48% 56%
Table 3. Test non-compliance results for tests generated from rules (RL) and tests generated from inverse rules (Inv) by PromptPex on different models. Higher test non-compliance (winner shown in blue) represents better tests.
@@
Figure 9. Average % Test Non-Compliance of tests generated by PromptPex and Baseline.
@@
Figure 10. Average % Test Non-Compliance of the tests generated for rules and inverse rules by PromptPex.
Figure 11. Number of valid tests generated by PromptPex.
Figure 12. Number of grounded rules generated by PromptPex.
4.7. How useful are rules for generating better tests? (RQ2)
We saw that PromptPex generates better tests using its approach, but we also evaluate how important is the fact that we generate tests for specific rules and inverse rules. As discussed in Section 3.3.2, we extract OR from the PUT. The OR is made up of rules which are the constraints over the output. We create inverse rules from these rules and use the generated rules to create tests. This approach allows us to compare the non-compliance of tests generated from the OR and from the inverse OR. Table 3 shows the percentage of their test non-compliance for different prompts on each model. The tests generated from the inverse rules on average generated 8.5% more non-compliant tests as shown in Figure 10 demonstrating the role of the OR and the rules whose inverse created these tests. Notably the tests generated for the inverse rules were able to consistently generate better tests even for a large model like gpt-4o-mini which otherwise is challenging even for the tests directly generated from PromptPex rules.
@@ 4.7. How useful are rules for generating better tests? (RQ2)
4.8. Does having an input specification help generate valid tests? (RQ3)
PromptPex uses IS to generate tests that meet the constraints on the input as described in PUT. In Figure 12, we show the total number of PromptPex generated tests for each benchmark as well as the number of tests considered valid by our validation analysis. We observe that a significant fraction of all generated tests are deemed valid, with the shakespeare prompt being an exception due to the input validator considering non-shakespearean input as invalid [^5] Writing prompts that process other prompts is challenging. In particular, our input validator in this case was confused and thought the input should be shakespearean text as well.. We also compare (not shown) the percentage of compliant valid tests with all the tests generated by PromptPex. We observed that 73% of all the non-compliant tests are valid tests. Non-complaint valid tests are valuable as they act as instant feedback and must be fixed.
4.9. Can automatically generated tests help determine if a model is suitable for a prompt? (RQ4)
@@
Prompt developers have many AI model options to choose from for a given application. They need to understand how effective a model is for their specific prompt. Because PromptPex automatically generates tests, we seek to understand if the automatically generated tests can appropriately distinguish model effectiveness for a given prompt.
@@
Figure 13 shows the PromptPex test % non-compliance for the models under test across the benchmarks. It illustrates that for a given prompt, tests generated by PromptPex strongly differentiates the capabilities of the models and can be used to determine model suitability. The more capable models (like gpt-4o-mini) have lower rates of non-compliance while the smaller models (like llama3.2:1b) have high rates. At the same time, some of the smaller models are highly capable for a given prompt. For example, both the gemma2:9b and quen2.5:3b models are effective for the speech-tag benchmark.
Figure 13. % Test Non-Compliance of tests generated by PromptPex.
4.10. Groundedness and Spec Agreement
It is important that the generated OR are both grounded in the PUT and cover all the important constraints expressed in the PUT.
@@ 4.10. Groundedness and Spec Agreement
In Figure 12, we show how many rules were generated for each benchmark and whether those rules were grounded, according to our groundedness evaluation. We found that on an average 89% rules from OR were grounded in the respective prompt from the benchmark with the classify prompt being an exception. The classify prompt takes input as a news article and classifies it into a given category. The non-grounded rules are generated by the model to compensate for the under-specification in the prompt. The classify prompt does not cover the cases when the output can be multiple categories or should the output just be the name of category without any explanation.
We observe high scores for spec agreement for all the prompts in our benchmark (96.8%) with the shakespeare prompt being an exception, without it the prompts achieved the spec agreement score of 99.9%.
@@
High groundedness and spec agreement score ensures that the OR generated is neither over or under extracted respectively and the tests generated by PromptPex are focused on testing the constraints in the PUT.
@@ 5. Discussion
In this section, we consider concrete examples from our benchmarks to better understand how PromptPex can help with the development of prompts. PromptPex not only generates tests but also generates specifications that are used for generating tests and have value on their own.
@@ 5.1. Input Specification
IS, which represents the constraints over the input, can be used to understand what the model expects as input for a particular prompt. This can be extremely useful when the input is complex or not obvious when reading the prompt. The developer can ensure the input spec remains the same with changes to the prompt that do not intend to change the input. It can also help in understanding if the input is under-specified in the prompt.
For example, for the elements prompt (Appendix B.6), the IS states that the input text can be of any length and can be in any language. This means that the model is expecting some implicit understanding of the input, as these constraints were not mentioned in the prompt itself, which only stated that the input is text.
In the art prompt (Appendix B.8), the description of the input in the prompt is confusing. The prompt takes user descriptions as input and converts them into prompts for generating art. The prompt talks about the input description and the generated prompt in a similar way, which is even confusing to read. It is not straightforward to separate the specification for the input and output. For example, ensuring each description does not exceed 80 words and is crafted in a single paragraph, the description refers to the input user description, but “crafted” can also mean it is referring to the generated prompt, which must be in a single paragraph. Such ambiguities are also reflected in the generated input specification, where an output constraint about the generated prompt being in English is present in IS. The prompt writer can take a look at the IS and update their prompt to make it clearer and without ambiguities.
Similarly, in the classify prompt (Appendix B.7), the prompt redefines the notion of a news article by stating that a news article can be classified into the following categories, which is reflected in the IS. If this was the actual intent, then it is okay; otherwise, seeing the IS provided the useful feedback to change the way a news article is defined in the prompt to make it more generic.
IS not only helps in generating tests that are valid but also helps the prompt developer understand if the input is properly specified in the prompt and if there is some ambiguity or errors around it.
@@ 5.2. Output Specification
The same can be said about the OR; it can also be used to understand what the model thinks about the output that can be generated. It can be used to figure out any under-specification or errors in the output constraints defined in the prompt.
For example, consider the elements prompt (Appendix B.6), where the values for multiple labels are extracted from the input. The OR states a rule that the labels must still be present even when there is no value that can be extracted for them. This case is not defined in the prompt, but it is a decision the model needs to take to handle such cases, thus pointing out an under-specification.
Similarly, for the art prompt (Appendix B.8), as we saw above while observing the IS, the OR is also getting confused, and there is a rule about only generating a description of length 80 words, which must be applied to the input. A prompt developer, after seeing the OR and IS, will be able to clarify the input and output in the prompt such that they generate better and clearer specifications, which also match the intent of the developer.
For the classify prompt, the OR adds a rule about not outputting anything other than the name of the category, pointing towards an under-specification, i.e., does it want the name of the category with or without any explanation, etc.
From these examples, it is clear that the specifications help in revising the prompt to improve any under-specification over the input and output. They can also be used to easily identify conflicting rules, which we did not see because the prompts we used were well-formed. The specifications can also be used to highlight bugs, like the inconsistencies found in the art prompt.
@@ 5.3. Inverse Rules
When we create tests from the rules in OR, we generate tests aimed at ensuring they follow the rule. Such tests ensure that the output is able to comply with the rules, and if it fails, they can help expose situations when there are conflicting rules causing the failure or when the model is not capable of generating output compliant with the rules. For a fairly capable model and a well-written prompt, these tests are expected to pass.
Since inverse rules contradict the rules themselves, they attempt to push the model to generate output that does not comply with the rules. We have shown an example of a test generated from the inverse rule in Figure 7, where the generated test is created such that it can potentially confuse the model and output multiple tags for the same word. These tests also have limitations as they are bounded by the input specification, which limits how challenging they can be. For example, for the speech tag prompt (Appendix B.1), in cases when the output can be Unknown or Can’t Answer, the test might contain a made-up word, a number, or an empty string. However, all of these are outside the input domain, which forces the test generator to create valid tests that are challenging.
@@ 5.4. Comparison with Baseline Tests
We examined the differences between the tests generated by PromptPex and those from the baseline to highlight some distinctive properties found in the tests created by PromptPex.
5.4.1. Creativity
@@
Tests generated by PromptPex are more creative than those produced by the baseline while still remaining within the input domain. For instance, in the speech tag prompt (Appendix B.1), PromptPex used non-existent words in tests like sentence: The truth was uncertain, shenative., word: shenative and sentence: The xylophone zxylophone harmonizes., word: zxylophone while the most creative attempt from the baseline was the use of 12 as a word, which lies outside the input domain. For the classify prompt (Appendix B.7), tests generated by PromptPex could be classified into multiple categories, like Google announces groundbreaking quantum computing progress, which fits both tech and business categories, unlike the baseline, which did not create ambiguous tests. In the extract name prompt (Appendix B.5), PromptPex produced many more tests with less common or imaginary model names such as ReinforceNet and AdvancedDL in varying contexts than the baseline did. For the sentence rewrite prompt (Appendix B.4), PromptPex addressed complex subject matter, such as the proliferation of digital technologies and integration of quantum computing, while the baseline focused on routine subjects. Lastly, in the Shakespeare prompt (Appendix B.3), PromptPex explored modern scenarios like Compose a modern dialogue about picking a TV show to watch and Write an excuse note for not doing my chores whereas the baseline was confined to traditional themes.
5.4.2. Complexity
PromptPex also produced more complex tests compared to the baseline. In the speech tag prompt (Appendix B.1), it used more complex words, for example, sentence: She spoke eloquently although about nothing in particular., word: although. For the classify prompt (Appendix B.7), PromptPex generated tests with varied descriptions, unlike the consistent tests from the baseline. In the sentence rewrite prompt (Appendix B.4), PromptPex used complex language and demonstrated conceptual depth with tests like juxtaposition of chaos and order in contemporary art, while baseline tests remained more literal. In the Shakespeare prompt (Appendix B.3), PromptPex introduced complexity by creating scenarios involving layered themes and character interactions, examples being Describe a character’s internal conflict in a scene and Write a dramatic scene about a kingdom’s downfall.

### v2

@@ 4.1.3. Benchmarks
We selected a diverse set of prompts from publicly available sources, focusing on those that are within the scope of PromptPex. Currently, we support prompts that can accept only a single input. Although this single input can be interpreted by the model to be made up of multiple components, this differentiation is not explicitly present, for example, a single string as input representing a sentence and a word for the speech tag prompt. This limitation makes it unsuitable for prompts requiring multiple embedded inputs. We also only support prompts where the output is independent of the previous outputs, making prompts describing multi-turn conversations for tasks out of scope for PromptPex.
4.1.4. Evaluation Procedure
@@
We accessed gpt-5_2025-08-07 and o4-mini_2025-04-16 through APIs. We kept the temperature 1.0 across all the requests. We used Ollama (20) for the local models, gpt-oss, gemma2:9b, qwen2.5:3b, llama3.2:1b. We ran all the experiments and hosted the local models on a virtual machine hosted in the cloud. Our system runs on an AMD EPYC 7V13 processor with 220 GB of RAM, 1 TB of SSD, and 4 NVIDIA A100 GPUs with a total of 80 GB of dedicated memory. The VM is running Ubuntu 24.04.3 LTS. We ran each test once per prompt per model.
4.2. Evaluation Results
Q1: Do specifications generate better tests?
@@ 4.2. Evaluation Results
Our hypothesis is that explicitly extracting output rules and using them to generate tests is more effective at generating tests that will cause non-compliance. But does our approach provide any benefits over baseline? To answer this question, we compare the non-compliance of tests generated by PromptPex and our baseline using 4 different models-under-test and averaged the scores across all the benchmarks. Figure 10 shows the average non-compliance across the benchmarks for PromptPex and baseline for each of the different models. More capable models will have lower non-compliance scores, as we see with the gpt-oss results when comparing against llama3:2.1b. Comparing against baseline, we see that the PromptPex tests have higher non-compliance than baseline tests for every model-under-test, but is particularly more effective for gpt-oss, where the baseline model has a harder time generating tests that result in non-compliant output.
@@
Figure 9. Average % Test Non-Compliance of tests generated by PromptPex and Baseline. Higher is better.
@@
Figure 10. Average % Test Non-Compliance of the tests generated for rules and inverse rules by PromptPex. Higher is better.
Q2: How useful are inverted rules in generating challenging tests?
Our hypothesis is that generating tests from inverted rules will create tests that are more likely to cause non-compliance. Because PromptPex generates tests from both output rules and inverted output rules, we can directly compare the non-compliance of both sets of tests to determine if inverting the rules is effective. Figure 10 compares the non-compliance of tests generated from rules and inverted rules for each of the models-under-test. The chart shows that tests from inverse rules result in greater non-compliance for all models-under-test. Furthermore, the best model-under-test, gpt-oss, shows the largest increase in non-compliance using tests from inverted rules.
Figure 11. Comprehensive comparison of test non-compliance rates between PromptPex tests (from rules and inverted rules) and baseline tests across all benchmarks for the gpt-oss model. Higher is better.
Figure 11 shows the per-benchmark results comparing baseline, rule-based, and inverted rule-based tests for the gpt-oss model. The figure illustrates that test compliance varies widely across the benchmarks but PromptPex tests (and especially invert-rule tests) result in higher non-compliance across the entire benchmark set.
@@
Q3: Does having an input specification help generate valid tests?
PromptPex extracts an input specification from the prompt so that it can generate tests that meet the defined constraints. Our goal is to ensure that the tests generated are valid according to that extracted specification.
Figure 12. Number of valid tests by PromptPex.
In Figure 12, we show the total number of PromptPex generated tests for each benchmark as well as the number of tests considered valid by our validation analysis. We note that almost every generated test is deemed valid based on the extracted input specification.
Q4: Can generated tests help determine if a model is suitable for a prompt?
Figure 13. % Test Non-Compliance of tests by PromptPex.
@@
Prompt developers have many AI model options to choose from for a given application and so they need to understand how effective a model is for their specific prompt. Because PromptPex automatically generates tests, we seek to understand if the automatically generated tests can appropriately distinguish the effectiveness of the model for a given prompt.
@@
Figure 13 shows the PromptPex test % non-compliance for the models under test across the benchmarks. It illustrates that for a given prompt, tests generated by PromptPex strongly differentiates the capabilities of the models and can be used to determine model suitability. The more capable models (like gpt-oss) have lower rates of non-compliance while the smaller models (like llama3.2:1b) have high rates. At the same time, some of the smaller models are highly capable for a given prompt. For example, both the gemma2:9b and qwen2.5:3b models are effective for the speech-tag benchmark.
@@
Groundedness and Spec Agreement It is important that the generated OR are both grounded in the PUT and cover all the important constraints expressed in the PUT. Using our groundedness evaluations, we find how many rules were generated for each benchmark and whether those rules were grounded. We found that on an average 89% rules from OR were grounded in the respective prompt from the benchmark with the classify prompt being an exception. The classify prompt takes input as a news article and classifies it into a given category. The non-grounded rules are generated by the model to compensate for the under-specification in the prompt. The classify prompt does not cover the cases when the output can be multiple categories or should the output just be the name of category without any explanation.
@@
We observe high scores for spec agreement for all the prompts (96.8%) with the shakespeare prompt being an exception, without it the prompts achieved the spec agreement score of 99.9%. High groundedness and spec agreement score ensures that the OR generated is neither over or under extracted respectively and the tests generated by PromptPex are focused on testing the PUT constraints.
@@ 5. Discussion
In this section, we consider concrete examples to better understand how PromptPex can help with the development of prompts.
@@ 5.1. Input Specification
IS can be used to understand what the model expects as input for a particular prompt. This can be useful when the input is complex or not obvious when reading the prompt. The developer can ensure the input spec remains the same with changes to the prompt that do not intend to change the input. It can also help in understanding if the input is under-specified in the prompt.
For example, for the elements prompt, the IS states that the input text can be of any length and can be in any language. This means that the model expects some implicit understanding of the input, as these constraints were not mentioned in the prompt itself, which only stated that the input is text. In the art prompt, the description of the input in the prompt is confusing. The prompt takes user descriptions as input and converts them into prompts for generating art. The prompt talks about the input description and the generated prompt in a similar way, which is even confusing to read. It is not straightforward to separate the specification for the input and output. For example, ensuring each description does not exceed 80 words and is crafted in a single paragraph, the description refers to the input user description, but “crafted” can also mean it is referring to the generated prompt, which must be in a single paragraph. Such ambiguities are also reflected in the generated IS, where an output constraint about the generated prompt being in English is present in IS. The prompt writer can take a look at the IS and update their prompt to make it clearer and without ambiguities. Similarly, in the classify prompt, the prompt redefines the notion of a news article by stating that a news article can be classified into the following categories, which is reflected in the IS. If this was the actual intent, then it is okay; otherwise, seeing the IS provided useful feedback to change the way a news article is defined in the prompt to make it more generic. IS not only helps in generating tests that are valid but also helps developers spot and fix ambiguities in input specifications.
@@ 5.2. Output Specification
OR can also be used to understand what the model thinks about the output that can be generated. It can be used to figure out any under-specification or errors in the output constraints defined in the prompt. For example, consider the elements prompt, where the values for multiple labels are extracted from the input. The OR states a rule that the labels must still be present even when there is no value that can be extracted for them. This case is not defined in the prompt, but it is a decision the model needs to take to handle such cases, thus pointing out an under-specification. Similarly, for the art prompt, as we saw above while observing the IS, the OR is also getting confused, and there is a rule about only generating a description of length 80 words, which must be applied to the input. A prompt developer, after seeing the OR and IS, will be able to clarify the input and output in the prompt such that they generate better and clearer specifications, which also match the intent of the developer. For the classify prompt, the OR adds a rule about not outputting anything other than the name of the category, pointing towards an under-specification, i.e., does it want the name of the category with or without any explanation, etc. Specifications help refine prompts by addressing under-specification in inputs and outputs. They also aid in detecting conflicting rules (though none were found in our well-formed prompts) and identifying bugs, such as inconsistencies in the art prompt.
@@ 5.3. Inverse Rules
Since inverse rules contradict the rules themselves, they attempt to push the model to generate output that does not comply with the rules. We have shown an example in Figure 7, where the generated test is created such that it can potentially confuse the model and output multiple tags for the same word. These tests also have limitations as they are bounded by the IS, which limits how challenging they can be. For example, for the speech tag prompt, in cases when the output can be Unknown or Can’t Answer, the test might contain a made-up word, a number, or an empty string. However, these are outside the input domain, which forces the test generator to create valid tests that are challenging.
@@ 5.4. Comparison with Baseline Tests
Tests generated by PromptPex are more creative and complex than those produced by the baseline while still remaining within the input domain. In the speech tag prompt, PromptPex used non-existent words in tests like sentence: The truth was uncertain, shenative., word: shenative and sentence: The xylophone zxylophone harmonizes., word: zxylophone while the most creative attempt from the baseline was the use of 12 as a word, which lies outside the input domain. For the classify prompt, tests generated by PromptPex could be classified into multiple categories, like Google announces groundbreaking quantum computing progress, which fits both tech and business categories, unlike the baseline, which did not create ambiguous tests. In the extract name prompt, PromptPex produced many more tests with less common or imaginary model names such as ReinforceNet and AdvancedDL in varying contexts than the baseline did. For the sentence rewrite prompt, PromptPex addressed complex subject matter, such as the proliferation of digital technologies and integration of quantum computing, while the baseline focused on routine subjects. Lastly, in the Shakespeare prompt, PromptPex explored modern scenarios like Compose a modern dialogue about picking a TV show to watch and Write an excuse note for not doing my chores whereas the baseline was confined to traditional themes. PromptPex introduced complexity by creating scenarios involving layered themes and character interactions, examples being Describe a character’s internal conflict in a scene and Write a dramatic scene about a kingdom’s downfall.
5.5. Comparison with Traditional Approaches
Another point of comparison for prompt test generation is specification-testing in traditional software, such as property-based testing. In this setting, tests are produced by (i) an input generator that samples valid inputs from the extracted input specification (IS) and (ii) an executable oracle derived from the extracted output rules (OR) that maps an input to an expected output (or checks required output properties). To evaluate our prompt-based test generation against this more symbolic alternative, we attempted to compile the extracted specifications into executable artifacts. We then used Hypothesis (maciver2019hypothesis) framework to sample up to 100 random inputs per prompt from the input_generator and recorded the resulting input–expected-output pairs as a conventional regression suite. This symbolic baseline is appealing because it enables regression tests without relying on an LLM judge at evaluation time. However, it breaks down for prompts with open-ended natural language outputs where there is no single reference answer. For a prompt like rewrite this paragraph more clearly while preserving meaning, an executable oracle cannot decide correctness and typically degenerates into shallow proxy checks (e.g., preserving named entities or forbidding new numbers), which are brittle.Moreover, randomized generation tends to explore the typical distribution induced by the strategy, whereas PromptTest’s rule and inverse-rule guided generation deliberately targets borderline-but-valid cases designed to stress specific constraints.
6. Related Work
Existing research has highlighted the need for prompt testing (Parnin et al., 2023; Ma et al., 2024) but has focused on using unit tests to assist model migration for a prompt (Dixit et al., 2024) or as a regression test suite to check future modifications to the prompt (Shankar et al., 2024). PromptPex automatically generates tests using the extracted specification from the prompt, and these tests can be used for model migration or as a regression test suite. PromptPex was informed by work on Program Exploration (Pex (Tillmann and de Halleux, 2008)), which automatically generates unit tests for .Net applications using dynamic symbolic execution.
Prompt Testing Currently prompt testing is mostly done manually or through benchmarks to evaluate a prompt’s performance on specific tasks like classification (authors, 2023). Model migration and regression also expect the user to provide a manually created unit test suite. Several prompt development platforms offer support for writing unit tests for prompts (8; 17; 11; 2; 24). Promptfoo (2) allows for writing assertions over generated output and SPADE (Shankar et al., 2024) automatically generates similar assertions over the output for checking regression but none of these platforms automatically generate unit tests.
Prompt Fuzzing Prompt fuzzing also addresses similar problems as PromptPex by generating tests to highlight flaws in the prompt. The key difference is that fuzzing also checks the prompt’s ability to handle ill-formed inputs, whereas PromptPex focuses on creating valid inputs. Even when it generates invalid tests, prompt fuzzing is closely related to our work as it involves generating tests for prompts. Currently, the primary focus of most prompt fuzzing efforts is the creation of adversarial inputs for red teaming (Yu et al., 2024; Yao et al., 2024; Gong et al., 2024; Yan et al., 2024; Yu et al., 2023). The automatic test generation in promptfoo (18) is also focused on generating malicious prompts. While exploring the impact of malicious inputs is valuable, PromptPex has a broader focus on generating inputs that help evaluate the robustness of the combination of the PUT and model under test.
Prompt Optimization Prompt optimization involves creating variants that perform better on a specific model or are more cost-effective. Although prompt optimization is not directly related to our work, most techniques for it require an equivalence checking test suite. Often, this suite consists of a set of manually labeled examples, either from the user or from common benchmarks (Pryzant et al., 2023; Chen et al., 2023; Sun et al., 2023; Khattab et al., 2023). Some methods use these datasets as seed examples and generate additional examples from them (Agarwal et al., 2024). The generation in such cases is more comparable to the baseline than to PromptPex.
Prompt Specification In PromptPex, we extract input and output specifications to serve multiple purposes, from understanding the prompt to generating input or output validators and unit tests. Stoica et al. (Stoica et al., 2024) have also highlighted the importance of specifications in prompt engineering to make it as reliable as traditional software engineering. Although we did not find any work on extracting and using specifications for generating tests, research exists on using specification for input and output validation and output generation. Sharma et al. (Sharma et al., 2024a) define input specifications for a VLLM prompt using a declarative meta-language, SPML (Sharma et al., 2024b). Amazon Bedrock Guardrails (23) allows defining policies over the output in a natural language abstraction over formal logic using declarative variables with natural language description. Structured or constrained decoding (10) generates output that follows a given domain-specific grammar, which is equivalent to an output structure specification.